A physical therapist wants to determine the difference in the proportion of men and women who participate in regular 15 physical activity what sample size should be obtained if she wishes the estimate to be within five percentage points with 90% confidence assuming that

a) she uses the estimate of 21.9% male and 19.3% female from a previous year?

b) she does not use any prior estimates?
Answer #1

Solution:- Given E = 0.05, 90% Confidence interval for Z = 1.645

formula n = (Z/E)^2*(p1q1+p2q2)

a) Given that p1 = 0.219,q1 = 0.781,p2 = 0.193,q2 = 0.807

n = (1.645/0.05)^2*((0.219*0.781)+(0.193*0.807))
= 353.7207

n = 354

b)no prior p1 = 0.5, q1 = 0.5, p2 = 0.5, q2 = 0.5

n = (1.645/0.05)^2*((0.5*0.5)+(0.5*0.5))
= 541.205

n = 542